The sum of two number is 90. If greater number is twice more than smaller number by 15. Find the number?
Answer by checkley77(12844)

X+Y=90
X=2Y+15
(2Y+15)+Y=90
3Y=90-15
3Y=75
Y=75/3
Y=25 ANS FOR THE SMALLER NUMBER.
X+25=90
X=90-25
X=65 ANS FOR THE LARGER NUMBER
65=2*25+15
65=50+15
65=65